I don't really know what the subects called but i don't know how to do equations with more than one exponent like 3p+7-(-3)=p+(-2) HELP! me please!
Answer by stanbon(75887)

3p+7-(-3)=p+(-2)
3p+7+3=p-2
3p+10=p-2
2p=-8
p=-4
Cheers,
Stan H.